OpenStudy (anonymous):

Solve the following system of equations using the substitution method. Enter your answers as (x,y) pairs, separated by commas. If the system is inconsistent, type INCONSISTENT. If the system is dependent, type DEPENDENT. x+3y=7 xy=4

OpenStudy (anonymous):

I got a sq root of y and i cant figure out the right equation at all

OpenStudy (anonymous):

(4,1)

OpenStudy (anonymous):

oops, no,, wait

OpenStudy (anonymous):

haha

OpenStudy (kinggeorge):

Well, \( x = 7-3y \) so \[ (7-3y)y = -3y^2 +7y = 4 \]so\[(-y+1)(3y-4)=0 \longrightarrow y = 1, \quad y = {4 \over 3}\]Testing these possibilities out, we find \(x = 4, \quad x = 3 \). Thus, our final solutions are \( (4, 1), (3, {4 \over 3}) \) Also, this is a dependent system since it has a solution.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (kinggeorge):

your welcome